8. Use the differential to approximate each quantity. (a) 0.99 (b) 145 (c) 20.01
Answer #1

8 a set fra)=ra f(a)= 2x Then by the detination flaton) - fea) = f(2) 42 or flatan)= feng + t(a) an 0 have to find the val(C) Let tead=ee f(a)=.ee row by the debination fraroa) f(a) = f(a) dx a f(a+09) = f(a) + f(926x 0.01 e we Shall find the v